# Exploring the Potential of Large Language Models in Simulink-Stateflow Mutant Generation

### v1

@@ Abstract.
$c$s: Software and its engineering Software verification and validationccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Flowcharts
@@ 1. Introduction
Mutation analysis (Budd, 1980) is used to assess the adequacy of test suites by inserting artificial faults into a program’s source code using mutation operators. These operators imitate developers’ common mistakes, producing modified program versions, known as mutants. The effectiveness of a test suite is then measured by its ability to detect behavioral differences between the mutants and the original program. A mutant is considered detected if at least one test case produces a different output than the original model; otherwise, it is considered alive. The mutation score, i.e., the percentage of mutants detected out of the total number of mutants, is the metric that determines the test adequacy. Despite its advantages, mutation testing has drawbacks, such as the generation of duplicated and equivalent mutants, high computational costs, and scalability issues.
@@
Conventional mutation analysis methods rely on rule-based mutation operators (Jia and Harman, 2010; Just, 2014), which make simple syntactic changes throughout the code at all possible locations. While effective at producing a large number of mutants, many are redundant or equivalent (i.e., mutants that exhibit no behavioral difference or are semantically identical to the original code), offering no additional value yet requiring increasing computational cost. To address this issue, recent studies (Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022) have explored using machine learning to generate mutants from historical project data and identified patterns. However, these techniques remain limited by the availability and quality of training data (Ojdanic et al., 2023), often leading to unrealistic or syntactically incorrect mutants that fail to accurately depict real-world faults. An alternative to this is to manually create mutants, but it is expensive and requires substantial domain expertise. Consequently, recent efforts have resorted to using Large Language Models (LLMs) for mutant generation (Tip et al., 2025b; Degiovanni and Papadakis, 2022), leveraging their extensive training on large-scale code data and their ability to reason over semantic and structural patterns. This enables LLMs to produce mutants that go beyond simple syntactic modifications, potentially generating more realistic and diverse mutants. Nevertheless, the effectiveness of LLM-generated mutants remains underexplored and requires explicit assessment.
@@
While the aforementioned challenges are well recognized in source-code mutation, they become even more pronounced in the context of Simulink-Stateflow models (MathWorks, 2023), which are widely used in developing embedded systems. Unlike traditional source code, Simulink-Stateflow models integrate continuous-time dynamics with discrete-event behaviors, requiring specialized mutation strategies to maintain syntactic validity, adhere to modeling standards, and reflect real-world faults. Simply altering syntax is often insufficient, as it may produce mutants that either violate model semantics or fail to mimic faults developers make in practice. Generating meaningful mutants, i.e., mutants that are syntactically valid, semantically consistent, and representative of realistic faults that could affect system behavior, is particularly difficult due to the unique semantics and hierarchical organization of Simulink-Stateflow models, including the connections between states, transitions, and actions triggered by events. Furthermore, as with source-code mutation, it is common to generate redundant and duplicated mutants (Fernandes et al., 2017), which increases the computational cost of current approaches due to the long execution times of Simulink-Stateflow models. Therefore, it is crucial to develop mutation testing approaches that consider the domain-specific attributes of Simulink-Stateflow models to produce high-quality, meaningful mutants.
@@
  • We extensively evaluate eight well-known LLMs, assessing their performance in generating Simulink-Stateflow mutants. We investigate how different mutation strategies, prompting strategies, and temperature settings affect the performance of LLMs.Our findings show that LLMs generate mutants up to 13 times faster than the baseline, with higher quality, greater diversity, and fewer duplicates or equivalents.
@@ 4.5.1. Effectiveness Metrics
Compilation Rate: Even when the generation of a proposed mutant is correct (i.e., the mutant parser in our approach succeeds in generating the Simulink-Stateflow model), not all generated mutants are guaranteed to compile. This metric measures the proportion of mutants that successfully compile after generation.
@@ 8. Conclusion
LLM-based mutant generation offers a promising new direction for mutation analysis for Simulink-Stateflow models, which are widely used in Cyber-Physical Systems (CPS) development. However,to the best of our knowledge their effectiveness across mutation strategies, prompting styles, and temperature settings had not been thoroughly studied in this domain. We present an automated pipeline that transforms Stateflow models into JSON and uses LLMs to generate mutants. In our evaluation, we evaluated eight state-of-the-art models with varying characteristics across eight well-known metrics in mutation analysis. In a large-scale study of 38,400 mutants across four Simulink-Stateflow models, LLMs generated mutants up to 13x faster than the baseline, with fewer duplicates and equivalents and higher overall quality. We reveal that global mutation with few-shot prompting achieved the highest generability and compilability, while local mutation improved quality but increased duplicate and equivalent rates. Moreover, temperature strongly influenced performance, with low–medium values being most effective. Finally, Error analysis showed four main failure causes: 1) invalid references, 2) incorrect transitions, 3) undefined variables, and 4) syntax errors, highlighting the need for structural and semantic validation. Long story short, LLMs showed to be effective and efficient at Simulink-Stateflow mutant generation.
@@ Acknowledgments
Pablo Valle is supported by the Pre-doctoral Program for the Formation of Non-Doctoral Research Staff of the Education Department of the Basque Government (Grant n. PRE_2025_2_0252). Shaukat Ali is supported by the Co-tester project (No. 314544) funded by the Research Council of Norway.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM Third International Conference on AI Foundation Models and Software Engineering; April 12–13, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM Third International Conference on AI Foundation Models and Software Engineering (FORGE ’26), April 12–13, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793655.3793737isbn: 979-8-4007-2477-0/2026/04ccs: Software and its engineering Software verification and validationccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Flowcharts
@@ 1. Introduction
Mutation analysis (Budd, 1980) is an effective software testing method used to assess the adequacy of test suites by inserting artificial faults into a program’s source code using mutation operators. These operators imitate developers’ common mistakes, producing modified program versions, known as mutants. The effectiveness of a test suite is then measured by its ability to detect behavioral differences between the mutants and the original program. A mutant is considered detected if at least one test case produces a different output than the original model; otherwise, it is considered alive. The mutation score, i.e., the percentage of mutants detected out of the total number of mutants, is the metric that determines the test adequacy. Despite its advantages, mutation testing has drawbacks, such as the generation of duplicated and equivalent mutants, high computational costs, and scalability issues.
@@
Conventional mutation analysistesting methods rely on rule-based mutation operators (Jia and Harman, 2010; Just, 2014), which make simple syntactic changes throughout the code at all possible locations. While effective at producing a large number of mutants, many are redundant or equivalent (i.e., mutants that exhibit no behavioral difference or are semantically identical to the original code), offering no additional value yet requiring increasing computational cost. To address this issue, recent studies (Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022) have explored using machine learning to generate mutants from historical project data and identified patterns. However, these techniques remain limited by the availability and quality of training data (Ojdanic et al., 2023), often leading to unrealistic or syntactically incorrect mutants that fail to accurately depict real-world faults. An alternative to this is to manually create mutants, but it is expensive and requires substantial domain expertise. Consequently, recent efforts have resorted to usingthe use of Large Language Models (LLMs) for mutant generation (Tip et al., 2025b; Degiovanni and Papadakis, 2022), leveraging their extensive training on large-scale code data and their ability to reason over semantic and structural patterns. This enables LLMs to produce mutants that go beyond simple syntactic modifications, potentially generating more realistic and diverse mutants. Nevertheless, the effectiveness of LLM-generated mutants remains underexplored and requires explicit assessment.
@@
While the aforementioned challenges are well recognized in source-code mutation, they become even more pronounced in the context of Simulink-Stateflow models (MathWorks, 2023), which are widely used in developing embedded systems. Unlike traditional source code, Simulink-Stateflow models integrate continuous-time dynamics with discrete-event behaviors, requiring specialized mutation strategies to maintain syntactic validity, adhere to modeling standards, and reflect real-world faults scenarios. Simply altering syntax is often insufficient, as it may produce mutants that either violate model semantics or fail todo not mimic faults made by developers make in practice. Generating meaningful mutants, i.e., mutants that are syntactically valid, semantically consistent, and representative of realistic faults that could affect system behavior, is particularly difficult due to the unique semantics and hierarchical organization of Simulink-Stateflow models, including the connections between states, transitions, and actions triggered by events. Furthermore, as with source-code mutation, it is common to generate redundant and duplicated mutants (Fernandes et al., 2017), which increases the computational cost of current approaches due to the long execution times of Simulink-Stateflow models. Therefore, it is crucial to develop mutation testing approaches that consider the domain-specific attributes of Simulink-Stateflow models to produce high-quality, meaningful mutants.
@@
  • We extensively evaluate eight well-known LLMs, assessing their performance in generating Simulink-Stateflow mutants. WIn our evaluation, we investigate how different mutation strategies, prompting strategies, and temperature settings affect the performance of LLMs.Our findings show that LLMs generate mutants up to 13 times faster than the baseline, with higher quality, greater diversity, and fewer duplicates or equivalents.
@@ 4.5.1. Effectiveness Metrics
Compilation Rate: Even when the generation of a proposed mutant is correct (i.e., the mutant parser in our approach succeeds in generatingto generate the Simulink-Stateflow model), not all generated mutants are guaranteed to compile. This metric measures the proportion of mutants that successfully compile after generation.
@@ 8. Conclusion
LLM-based mutant generation offers a promising new direction for mutation analysis for Simulink-Stateflow models, which are widely used in Cyber-Physical Systems (CPS) development. However,to the best of our knowledge their effectiveness across mutation strategies, prompting styles, and temperature settings had not been thoroughly studied in this domain. We present an automated pipeline that transforms Stateflow models into JSON and uses LLMs to generate mutants. In our evaluation, we evaluated eight state-of-the-art models with varying characteristics across eight well-known metrics in mutation analysis. In a large-scale study of 38,400 mutants across four Smimulink-Stateflow models, LLMs generated mutants up to 13x faster than the baseline, with fewer duplicates and equivalents and higher overall quality. We reveal that global mutation with few-shot prompting achieved the highest generability and compilability, while local mutation improved quality but increased duplicate and equivalent rates. Moreover, temperature strongly influenced performance, with low–medium values being most effective. Finally, Error analysis showed four main failure causes: 1) invalid references, 2) incorrect transitions, 3) undefined variables, and 4) syntax errors, highlighting the need for structural and semantic validation. Long story short, LLMs showed to be effective and efficient at Simulink-Stateflow mutant generation.
@@ Acknowledgments
Pablo Valle and Aitor Arrieta are part of the Software and Systems Engineering research group of Mondragon Unibertsitatea (IT1519-22), supported by the Department of Education, Universities and Research of the Basque Country. Pablo Valle is supported by the Pre-doctoral Program for the Formation of Non-Doctoral Research Staff of the Education Department of the Basque Government (Grant n. PRE_2025_2_0252). Shaukat Ali is supported by the Co-tester project (No. 314544) funded by the Research Council of Norway.